Calculate the radius of 1st 2nd 3rd orbits of hydrogen and helium atom

Grade:10

2 Answers

Arun
4 years ago
Radius of bohr orbit i=0.529n^2/z
as for hydrogen z=1
1st orbit=0.529
2nd orbit n=2 and z=1 so =2.116
3rd orbit n=3 and z=1 so=4.775
Vikas TU
4 years ago
Bohr's Radius is given by 
rn = 0.529*n^2 /z Angestrum 
n = no. of energy level 
Z=1 
Radius of irst energy level = n=1 
r1 = 0.529* 1^2 /1 = 0.529 A
Raius of second energy level = n=2 
r2 = 0.529 * 4/1 = 2.116 A 
Radius of third energy level = n=3 
r3 = 0.529* 9/1 = 4.761 A